OpenStudy (anonymous):

hi there. let me ask you a question, please answer badly needed. In polar, there are different kinds of graphs how do i determine if it is cardioid, circle, limacons, flower or spiral?

Directrix (directrix):

I learned four or five of the basic polar graph equation formats and the type graph that went with them. Just as in Cartesian, I learned several formats so that I would recognize those. For example, y = x^2 is the parabola. In polar, r = sin( theta) is a circle. But, there's no memorizing all of them. I used a calculator to crank out some of the graphs. Other polars, I converted to Cartesian.

Directrix (directrix):

@jaja22 How are you learning the which equation goes to which graph? It just came to me that I misunderstood your question. Do you want to know how to convert a Cartesian equation to a Polar one or the other way around? I am speaking of the Algebra involved. Here is a link to a good resource on the topic: http://sites.csn.edu/istewart/mathweb/math127/polar_equ/polar_equ.htm

Directrix (directrix):

If you have a particular problem you'd like help in solving, please post it.

OpenStudy (anonymous):

ahm in our past lesson if the equation goes like this r=a+b cos theta the graph is limacon with a loop so it means we can picture out what graph could it be.

Directrix (directrix):

The cardioid and the dimpled limacon are special cases of the generic limacon as best I remember. I have a Catalog of Plane Curves. I will look later.

Directrix (directrix):

I do not think I have answered your question.

OpenStudy (anonymous):

it's ok, thanks anyway. you said cardioid and limacon are special cases. why?

Directrix (directrix):

Read the information at the link: http://mathworld.wolfram.com/Limacon.html

OpenStudy (anonymous):

@Directrix thank you for your help